OpenStudy (anonymous):

A car moving at a speed of 20 m/s, then accelerates uniformly at 1.8m/s^2 until it reaches a speed of 25m/s What distance does it travel during the time it is accelerating?

OpenStudy (anonymous):

btw, are you doing your homework? XD 25 m/s - 20 m/s = 5 m/s, 5 m/s / 1.8 m/s^2 = 2.7777778 s D = Vit + 1/2at^2 D = 20 m/s * 2.778 s + 1.8 m/s^2 * (2.778s)^2 D = 55.56m + 5.0004m D = 60.5604m

OpenStudy (anonymous):

0.16 m/s2 0.20 m/s2 0.14 m/s2 0.19 m/s2 0.17 m/s2 ? I'm sorry but tbh physics is harddddd

OpenStudy (anonymous):

???? it asks for the distance, why are the options in acceleration?

OpenStudy (anonymous):

wrong answers haha these are 58 m 50 m 54 m 66 m 62 m

OpenStudy (anonymous):

.-. closest answer is 62m

OpenStudy (anonymous):

Thank you XD